OpenStudy (anonymous):

f(x)=x^2+3x+4. determine the x value if y=2 and y=8

OpenStudy (solomonzelman):

x²+3x+4=2 x²+3x+2=0 (x+1)(x+2)=0 x = ?

OpenStudy (solomonzelman):

finish it.. I am going to Kroger now....

OpenStudy (solomonzelman):

For y=2 x²+3x+4=2 x²+3x+2=0 (x+1)(x+2)=0 x= -1, -2 For y=8 x²+3x+4=8 x²+3x-4=0 (x-1)(x+4)=0 x= 1, -4

OpenStudy (anonymous):

thank you vary much i get ti now